Express the function g ,defined by g(x)=1/√x+√x,as a composition of three function


1
Expert's answer
2022-04-19T04:44:09-0400

"f(x)=\sqrt{x},\;g(x)=\sqrt{x^2+x},\;h(x)=\frac{1}{x}."

"g(f(x))=\sqrt{(\sqrt{x})^2+\sqrt{x}}=\sqrt{x^2+x}."

"h(g(f(x)))=\frac{1}{g(f(x))}=\frac{1}{\sqrt{x^2+x}}."